The average time it takes a group of senior high school students to

complete a certain examination is 46.2 minutes. The standard deviation is

8 minutes. Assume that the variable is normally distributed.


What is the probability that a random selected college student will

complete the examination in less than 43 minutes?


1
Expert's answer
2022-05-06T12:36:56-0400

Suppose that "X" is a random variable that has a normal distribution with parameters "\mu=46.2" and "\sigma=8". The respective probability density function is: "p(x)=\frac{1}{\sigma\sqrt{2\pi}}e^{-\frac12\left(\frac{x-\mu}{\sigma}\right)^2}". The aim is to find "P(X\leq43)". We receive: "P(X\leq43)=\int_{-\infty}^{43}\frac{1}{\sigma\sqrt{2\pi}}e^{-\frac12\left(\frac{x-\mu}{\sigma}\right)^2}dx\approx0.3446" (it is rounded to "4" decimal places)

Answer: "0.3446" (it is rounded to "4" decimal places)